Main Logbooks page LST commissioning Shift checklists Safe state checklists Maintenance activities LST2-4 Commissioning Old logbooks
GENERAL CAMERA DRIVE ENERGY AMC
  LST1 Drive ELOG  ELOG logo
Message ID: 245     Entry time: Mon Apr 1 07:43:23 2019
Author: Drive Team- Ino&Thierry&Armand 
Type: Problem Fixed 
Category: Hardware 
Subject: LST Parking control 
Status:  
  • On Wenesday morning (27/03/19) :
    • With Holger and Robert (MPI) , many telescope "park in/park out" operations have been performed in order to refine the positionning in the acces tower and at the elevation locking bolt
      Holger and Robert applied many mechanical modifications on the acces tower to make the telescope park operations very smooth
      ==> Conclusive operations. 
ELOG V3.1.5-fc6679b